OpenStudy (anonymous):

f(x)=x^3-3x^2+4 find the zeros? I know what the answer is, I just don't know how to get it...could you explain in detail please?? and what is a duplicate root?

OpenStudy (amistre64):

factor, trial and error, etc. a duplicate root might suggest that the factored form has 2 of the same factors

OpenStudy (amistre64):

a rational roots method might bring about something close; and narrow down any trial and error

OpenStudy (amistre64):

\[\pm\frac{factors.of.last}{factors.of.first}\]

OpenStudy (amistre64):

since the first number is 1; it makes life alot easier :)

OpenStudy (amistre64):

so factors of 4 are: 1,2,4 plus or minus of course

OpenStudy (amistre64):

lets see if any of these makes a zero .. f(x) = x^3-3x^2+4 f(1)=1-3+4 = not 0 f(-1)=-1-3+4 = 0 ... -1 is a root f(2) = 8-12+4 = 0 ... 2 is a root f(-2) = -8-12+4 = not 0 f(4) = 64-48+4 = not 0 f(-4) = -64-48+4 = not 0

OpenStudy (amistre64):

so we know that (x+1)(x-2) can be divided out of it: and that we are missing another root, since its a 3degree poly. That means that one of these most likely is duplicated

OpenStudy (amistre64):

logic would says its the (x-2)

OpenStudy (anonymous):

thanks!! I get it now. :DD